DX 4D flattening of TS3 - 20180205

OVMS Telemetry is being saved in parallel to the collection of 4D data.

17:59
  • print, opt_int_mat_reduce2(30, track='20180203_215736')
  • remove mostly astigmatism
  • flat 175845
  • print, hexalign()

18:04 Both LUCI's On, Telescope at zenith, HBS on, LFBG@341, RFBG@210
  • comm4d.capture(1000, '20180205_180400_noise') - chiller ? (reduced power)
  • comm4d.capture(1000, '20180205_180600_noise') - chiller ?
  • comm4d.produce('20180205_180600_noise')
  • asma_opt_noise_new, '20180205_180600_noise')

18:18 John M. report the the LUCI cryo pumps are off (all 4 of them)

18:xx Both LUCI's Off
  • comm4d.capture(1000, '20180205_185100_noise') - chiller turning on
  • comm4d.capture(1000, '20180205_185200_noise') - chiller ON
  • comm4d.produce('20180205_185200_noise')

19:04 John M. Turn on LUCI2 cold head #1

19:32 LUCI2 Cold Head #1 On, others off
  • comm4d.capture(1000, '20180205_193200_noise') - chiller ON
  • comm4d.capture(1000, '20180205_193500_noise') - chiller ?
  • comm4d.capture(1000, '20180205_194300_noise') Improved Fringes by changing T/T (Rx, Ry) - chiller OFF
  • hexapod 0.95, -0.775, 2.54, 134.3, 341.3

19:47 LUCI1 both Cold Heads On, LUCI2 Cold Head #2 On, LUCI2 Cold Head #1 Off
  • comm4d.capture(1000, '20180205_194700_noise') - chiller OFF
  • comm4d.capture(1000, '20180205_195000_noise') - chiller OFF
  • comm4d.capture(1000, '20180205_195300_noise') - chiller ON

21:10 LBTI Off, LUCI1 both Cold Heads On, LUCI2 Cold Head #2 On, LUCI2 Cold Head #1 Off
  • comm4d.capture(1000, '20180205_201000_noise') - chiller OFF
  • comm4d.capture(1000, '20180205_201100_noise') Exp time 47 us (5) - chiller OFF
  • comm4d.capture(1000, '20180205_201300_noise') Exp time 31 us (0) - chiller OFF
  • comm4d.capture(1000, '20180205_201700_noise') Exp time 31 us (0) - chiller OFF

21:20 LBTI On, Both LUCIs On
  • comm4d.capture(1000, '20180205_204400_noise') Exp time 47 us (5) - chiller OFF
  • comm4d.capture(1000, '20180205_204700_noise') Exp time 31 us (5) - chiller OFF
  • Move Rx, Ry to reduce number of fringes
  • comm4d.capture(1000, '20180205_205000_noise') Exp time 31 us (5) - chiller OFF
  • comm4d.capture(1000, '20180205_205100_noise') Exp time 47 us (5) - chiller OFF

21:09 Install Guido's new 17 iteration ff matrix
  • cp /local/aomeas/adsec_calib/ff_matrix/20180204_210014/filtered/*.fits /home/aoeng/confcalib/calib/adsec/672a/data/.
  • adsc_stop/start

21:12 Load fewer modes
  • print, fsm_load_shape('..../default.sav', force_max=50)

21:34 Flatten with new FF matrix
  • print, opt_int_mat_reduce2(100, track='20180203_215736')
    • after 100m old RMS = 101 nm
    • flat 213438
  • print, opt_int_mat_reduce2(300, track='20180203_215736')
    • after 300m old RMS = 55 nm
    • flat 213749

21:52 Measure new Recon 301-500
  • print, id4_opt_intmat(301, 500) routine chooses amplitude
    • 20180205_215203
  • comm4d.produce('20180205_215203')
  • print, if_redux(/modal, '20180205_215736')
  • print, build_int_mat(['20180203_215736', '21080203_223240', '20180205_215203'])
    • before 150m RMS = 70 nm
  • print, opt_int_mat_reduce2(50, track='20180203_215736')
    • after 500m RMS = 52nm
    • flat 224648

22:47 Collect new 0-500 mode data
  • print, id4_opt_intmat(0, 200)
    • 224958 (max ptv tilt (surface) nm: 1321) (RMS tilt coif (surface) nm: 948)
    • 225511 (max ptv tilt (surface) nm: 1267) (RMS tilt coif (surface) nm: 928)
  • print, id4_opt_intmat(201, 350)
    • 225250 (max ptv tilt (surface) nm: 1341) (RMS tilt coif (surface) nm: 1013)
    • 225657 (max ptv tilt (surface) nm: 1296) (RMS tilt coif (surface) nm: 918)
  • print, id4_opt_intmat(351, 500)
    • 225357 (max ptv tilt (surface) nm: 1353) (RMS tilt coif (surface) nm: 972)
    • 225752 (max ptv tilt (surface) nm: 1163) (RMS tilt coif (surface) nm: 901)

20180208 do if_redux() and build_opt_intmat()
  • 500 mode recon in 20180205_225511

23:08 Set default.sav to ./20180205_500m.sav
  • See high current on act 376 (8:00 on mirror GUI)

23:11 DX Hexapod 0.98. -0.788. 2.54, 136.5, 346.0

23:11 Hand over to mountain for RR removal

Preliminary Conclusions from Vibration Testing with 4D and OVMS

J. Hill on behalf of many.....

a) The 4D optical Tip-Tilt data taken this morning (20180205) is 3x less noisy than it was on Saturday afternoon (20180203) with telescope in same configuration (telescope at zenith, with HBS on, all LUCI coldheads running) WHY? - there must be some alternate source of TT noise. The conclusion seems to be that the LUCI cold heads are not the dominant source of 4D TT noise even though there are some corresponding frequency peaks. (This is based on the 28 Hz sampling of the 4D wavefront data.) Does anyone have any ideas for the source of the alternate noise which is apparently below 20 Hz?

[switching gears here to the OVMS data rather than 4D data]

b) LUCI2 Coldhead #1 (the same location as the one that was replaced in Nov 2017) is responsible for essentially all of the DX TT PSD peaks (OVMS+) that are overtones of 1.2 Hz. These include the peaks at 18.0, 42.0, 58.8, 60.0 Hz. In some instances the amplitude of these narrow peaks (i.e. 42.0 Hz) were reduced by a factor of 1000 when the coldhead was switched off. Above 60 Hz, this coldhead is still 10x quieter than the one that it replaced (e.g. 61.2 Hz). At 60 Hz and below, the amplitudes are similar before and after the replacement.

c) The other 3 LUCI Coldheads do not contribute in any obvious way to the DX TT PSD spectrum. Their contribution is un-noticeable, although perhaps not completely un-measurable.

d) Today (20180205) there is a new peak in the DX TT PSD at 81 Hz which was NOT present on Friday (20180202). While comparable in amplitude to the 60 Hz peaks, this 81 Hz peak was not influenced by turning off the LUCI cold heads nor the LBTI cold head. This 81 Hz peak seems to have grown by 2 orders of magnitude in the first 3 hours of 20180204 UT. We see this peak sometimes in the past, but never (in 20-ish samples) more than 10% of its present amplitude (e.g. 12 UT on 20180121).

e) We also note a peak at 30.0 Hz in the SX Tip PSD which is not influenced by the LUCI cold heads.


-- %USERSIG{DougMiller - 2018-02-05}%

Comments

 
  • OVMS TT:
    ovms.mode2.20180205bothLUCIonExceptLUCI2ch1LBTIoff.jpg

  • OVMS TT:
    ovms.mode2.20180205bothLUCIonExceptLUCI2ch1.jpg

  • OVMS TT for Both LUCI On:
    ovms.mode2.20180205bothLUCIon.jpg

  • OVMS TT for LUCI2 ch1 only on:
    ovms.mode2.20180205LUCI2ch1ONLY.jpg

  • OVMS TT for all four LUCI cold heads off:
    ovms.mode2.20180205bothLUCIoff.jpg
I Attachment Action Size Date Who Comment
ovms.mode2.20180205LUCI2ch1ONLY.jpgjpg ovms.mode2.20180205LUCI2ch1ONLY.jpg manage 1 MB 09 Feb 2018 - 21:03 JohnHill OVMS TT for LUCI2 ch1 only on
ovms.mode2.20180205bothLUCIoff.jpgjpg ovms.mode2.20180205bothLUCIoff.jpg manage 1 MB 09 Feb 2018 - 21:04 JohnHill OVMS TT for all four LUCI cold heads off
ovms.mode2.20180205bothLUCIon.jpgjpg ovms.mode2.20180205bothLUCIon.jpg manage 1 MB 09 Feb 2018 - 21:02 JohnHill OVMS TT for Both LUCI On
ovms.mode2.20180205bothLUCIonExceptLUCI2ch1.jpgjpg ovms.mode2.20180205bothLUCIonExceptLUCI2ch1.jpg manage 1 MB 09 Feb 2018 - 21:01 JohnHill OVMS TT
ovms.mode2.20180205bothLUCIonExceptLUCI2ch1LBTIoff.jpgjpg ovms.mode2.20180205bothLUCIonExceptLUCI2ch1LBTIoff.jpg manage 1 MB 09 Feb 2018 - 21:00 JohnHill OVMS TT
Topic revision: r7 - 09 Feb 2018, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback